OpenStudy (anonymous):

3-sqrt(x)/9-x. how do you break down?

OpenStudy (unklerhaukus):

\[\frac{3-\sqrt x}{9-x}\] factorize the denominator use \[(a+b)(a-b)=a^2-ab+ba-b^2=a^2-b^2\] cancel common factors

OpenStudy (anonymous):

yes i know that but how do you that for sqrt?

OpenStudy (unklerhaukus):

use \(x=\sqrt x^2\)

OpenStudy (anonymous):

mmm so 3-sqet(x)=(x^2-3^9)?

OpenStudy (unklerhaukus):

your looking at the numerator (bottom bit of the fraction) you should be looking at the denominator ( the bottom bit of the fraction) \[9-x=(\cdots)(\cdots)\]

OpenStudy (anonymous):

9-x...... sqert(3)-(x)??

OpenStudy (unklerhaukus):

\[9-x=3^2-\sqrt x^2=(\cdots)(\cdots)\]

OpenStudy (anonymous):

(3-sq\[\left( 3-\sqrt{x} \right)\]

OpenStudy (anonymous):

(3-sq\[\left( 3-\sqrt{x} \right)\](3-sq\[\left( 3-\sqrt{x} \right)\]?

OpenStudy (anonymous):

\[\left( 3-\sqrt{x} \right)\]

OpenStudy (anonymous):

\left( 3-sqrt{x} \right) \left( 3-sqrt{x} \right)

OpenStudy (anonymous):

(3-sqrt(x)) (3-sqrt(x)

OpenStudy (unklerhaukus):

not quite\[a^2-b^2=(a+b)(a-b)\] so\[9-x=3^2-\sqrt x^2=\]

OpenStudy (anonymous):

3-sqrt(x)?

OpenStudy (anonymous):

hope its right

OpenStudy (unklerhaukus):

\[9-x=3^2-\sqrt x^2=(3+\sqrt x)(3-\sqrt x)\]

OpenStudy (unklerhaukus):

so\[\frac{3-\sqrt x}{9-x}=\frac{3-\sqrt x}{(3+\sqrt x)(3-{\sqrt x)}}\]\[\qquad\qquad =\frac{\cancel{3-\sqrt x}}{(3+\sqrt x)\cancel{(3-{\sqrt x)}}}\]

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

what about x^3+64/x+4

OpenStudy (anonymous):

I got (x+4)(x^2+4x+16)/(x+4)?

OpenStudy (unklerhaukus):

almost \[(a^3+b^3)=(a+b)(a^2-ab+b^2)\]

OpenStudy (unklerhaukus):

once you have fixed the sign error, cancel the common factor

OpenStudy (anonymous):

Thank you it was very helpful!!